OpenStudy (waheguru):

I Dont understand this question - Identify equal rational numbers in the list that follows 2/3 -3/2 -2/3 2 / 3 (a negative sign beside the line separating the 2 numbers) 3/2 (a negative sign beside the line separating the 2 numbers) 2/-3 3/-2 and 3/2

OpenStudy (waheguru):

rational numbers question

OpenStudy (anonymous):

yes i see that

OpenStudy (waheguru):

.

OpenStudy (anonymous):

what dont you understand?

OpenStudy (anonymous):

foil it out

OpenStudy (waheguru):

here is the question

OpenStudy (waheguru):

i dont understand what i am suppose to do

OpenStudy (anonymous):

Ahh...ok...now i see. A negative fraction is the same no matter where the minus sign is. So... \[\frac{-3}{2}=\frac{3}{-2}=-\frac{3}{2}\]

OpenStudy (waheguru):

what do i do i am soo confused

OpenStudy (anonymous):

find all the numbers that are the same given what I just mentioned above.

OpenStudy (waheguru):

is there a order i write them in?

OpenStudy (waheguru):

or can i just write them in any order but they have to be equal

OpenStudy (anonymous):

as far as i can tell there's no particular order they are looking for. as long as they are equal. you have two different fractions that are represented the way I have above...just so you know.

OpenStudy (waheguru):

so

OpenStudy (anonymous):

I gave you one of the two...what other fraction has the same forms?